Joe wants to build a rectangular shaped deck whose length will be three inches greater than its width. What will be the dimensions of the deck if the perimeter is to be 54 inches.
Answer by oberobic(2304)

P = perimeter
P = 54 in.
P = 2(L+W), where L=Length and W=Width
.
L = W+3
.
substitute
.
2(W+3) + 2W = 54
.
2W +6 +2W = 54
.
4W = 48
.
W = 12
.
L = W+3
L = 12+3
L = 15
.
Check the perimeter to be sure this solution is the answer.
.
2L = 2(15) = 30
2W = 2(12) = 24
30+24 = 54
Correct.
.
Answer: Length = 15 inches; Width = 12 inches.
.
Is this a deck for a doll house? Just wondering...(smile)